Selected data from Kamal Company's financial statements is as follows:

In $000's

2017

2016

Current liabilities

$ 5,130

$ 5,202

Long-term debt

1,419

1,322

Stockholders' equity

12,632

11,166

Interest and principal payments

300

400

Net income

2,430

2,280

Interest expense

196

172

Income taxes

1,837

1,524

Dividends paid

750

750


Kamal Company's times interest earned ratio for 2017

Select one:

A. shows an increase in the company's ability to make its interest payments.

B. increased, which indicates the company's creditors will be pleased.

C. indicates the company cannot meet its current year interest payments out of current year earnings.

D. decreased slightly, which indicates the company has about the same ability to pay interest on its debt.

Selected data from Kamal Company's financial statements is as follows:

In $000's

2017

2016

Current liabilities

$ 5,130

$ 5,202

Long-term debt

1,419

1,322

Stockholders' equity

12,632

11,166

Interest and principal payments

300

400

Net income

2,430

2,280

Interest expense

196

172

Income taxes

1,837

1,524

Dividends paid

750

750


Kamal Company's debt-to-equity ratio was 0.52 to 1 in 2017 and 0.58 to 1 in 2016.

Which of the following statements is true concerning Kamal?

Select one:

A. The company is improving its debt-to-equity ratio.

B. The company relied more on creditors for capital during 2017 than in 2016.

C. The company appears to be in a weaker position at the end of 2017 to finance capital expenditures from cash flow generated by operating activities.

D. The company has a smaller percentage of capital from owners at the end of 2017 than at the end of 2016.
Answer #1

Interest Earned Ratio= Earnings Before Income and Taxes/Interest

Interest Earned Ratio for 2016= (2280+172+1524)/172 =23.12

Interest Earned Ratio for 2016= (2430+196+1837)/196 =22.77

So the answer is D. decreased slightly, which indicates the company has about the same ability to pay interest on its debt..

Debt to equity ratio =Debt/Equity

Any Increase in the Debt to Equity Ratio means Company is more dependant on the Debt and vice versa. In this Case the devt to equity ratio has decreased as provided in the question which means owners capital has increased which has improved its ratio .

So the answer is A. The company is improving its debt-to-equity ratio.